# Pushing Toward the Simplex Vertices: A Simple Remedy for Code Collapse in Smoothed Vector Quantization

### v2

@@ 2.3 Regularization
$\mathcal{L}_{\text{reg}}$ $=\mathcal{L}_{\text{hard}}:=N^{-1}\sum_{i=1}^{N}\left(\beta\underbracket{\|\mathbf{z}_{i}-\mathtt{detach}(\mathbf{q}_{\iota(\mathbf{z}_{i})})\|^{2}}_{\text{Commitment Loss}}+\underbracket{\|\mathtt{detach}(\mathbf{z}_{i})-\mathbf{q}_{\iota(\mathbf{z}_{i})}\|^{2}}_{\text{Codebook Loss}}\right)$ (10)
@@
It should be noted, however, that the codebook loss in Equation 10 does not inherently prevent code collapse. Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates (Zhu et al.,, 2025). Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features Dhariwal et al., (2020). More recently, Zhu et al., (2025) proposed another remedy called SimVQ, which reparameterizes the codebook $Q$ as the product of a randomly frozen matrix $Q^{\prime}\in\mathbb{R}^{M\times D}$ and a learnable matrix $W\in\mathbb{R}^{D\times D}$ (i.e., $Q=Q^{\prime}W$). In this formulation, all codebook vectors share learnable parameters with one another ($\mathbf{q}_{m}=q^{\prime}_{m,1}\mathbf{w}_{1}+\cdots+q^{\prime}_{m,D}\mathbf{w}_{D}$), so updates to one vector propagate to the others.
@@ 4 Experiments
The proposed regularization for smoothed vector quantization was benchmarked on two tasks: discrete autoencoding (§4.1) and contrastive learning (§4.2). The Python code used for these experiments is available as supplementary material.
@@ 4.1 Discrete Autoencoding
The first benchmark assessed the proposed regularization in the context of discrete autoencoding on the ImageNet dataset (Deng et al.,, 2009). Input images were convolutionally encoded into latent feature maps, whose pixels were then quantized (Esser et al.,, 2021; Fifty et al.,, 2025). The decoder convolutional network reconstructed the input images from these quantized feature maps, and the entire model was trained to minimize the L2 reconstruction loss ($L_{\text{main}}$ in Equation 9). Further details about the network architecture and training setup are provided in Appendix B.1.
Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), three different combinations of feature map and codebook sizes were examined. In addition to the previously used settings of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), an additional configuration increased the channel dimensionality of the latter to $32$ ($H\times W\times C=64\times 64\times 32,M=8196$).
Each model was trained using four GPUs, and the proposed method computed the $K/4$-nearest neighbors of each simplex vertex per GPU (see §5.3 for further discussion). The number of neighbors was set to $K/4=8$ for the L2 distance and $K/4=1$ for the cross-entropy otherwise specified. [^3] The value of $K$ was upper-bounded at $8\times 4$ by available computational resources; With a maximum batch size of $64$, the total number of latent pixels was $64\times 64\times 64=8\times 4\times 8196$, allowing only $8\times 4$ neighbors per vertex of $\Delta^{8196}$. This implementation constraint is further discussed in §5.3. The weight $\beta$ on the commitment loss (Equation 10) in STE (including SimVQ) and rotational gradient estimation was set to 1.0, following Fifty et al., (2025) and Zhu et al., (2025).
@@
$\times16$$\times$32; 1024 $\times64$$\times$3; 8196 $\times64$$\times$32; 8196
Method $K/4$ Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS
STE
Euclid 4.5% 0.404 124.86 36.61 100.0% 0.167 7.25 402.84 1.7% 0.235 22.02 290.24
Cosine 3.0% 0.381 117.95 41.48 70.9% 0.197 13.97 348.31 2.8% 0.186 12.11 363.86
SimVQ 100.0% 0.340 87.26 71.33 100.0% 0.170 7.44 400.41 100.0% 0.148 3.97 436.29
RE
Euclid 3.1% 0.460 170.30 19.25 78.85% 0.171 10.21 377.56 0.5% 0.271 40.58 203.52
Cosine 2.8% 0.423 157.77 24.33 99.5% 0.194 14.67 344.17 4.4% 0.180 10.83 372.89
H-Gmb
PPL 100.0% 0.349 100.29 54.04 100.0% 0.189 19.07 321.16 100.0% 0.163 10.33 384.04
KNN-L2 8 52.1% 0.344 98.98 59.66 99.9% 0.222 25.00 280.97 23.6% 0.185 10.72 376.35
KNN-CE 1 100.0% 0.368 86.49 68.27 100.0% 0.226 14.27 342.64 100.0% 0.173 3.17 435.04
S-Gmb
PPL 55.0% 0.826 173.59 9.61 100.0% 0.183 10.28 377.79 48.4% 0.386 45.61 167.14
KNN-L2 8 96.3% 0.358 74.00 85.36 100.0% 0.201 10.17 373.83 100.0% 0.193 5.73 404.18
KNN-CE 1 100.0% 0.569 194.74 16.60 100.0% 0.233 15.41 329.78 100.0% 0.186 3.83 425.33
Softmax
PPL 100.0% 1.112 309.60 4.55 81.6% 0.701 37.99 196.72 99.8% 0.725 83.07 77.98
KNN-L2 1 100.0% 0.358 76.29 81.61 100.0% 0.224 12.08 354.36 100.0% 0.387 32.20 210.61
2 100.0% 0.366 81.69 71.75 100.0% 0.226 12.04 348.48 100.0% 0.175 3.62 427.42
4 100.0% 0.379 87.05 62.37 100.0% 0.205 9.01 383.56 100.0% 0.196 5.64 403.06
8 99.9% 0.343 73.72 88.53 100.0% 0.199 12.64 361.23 100.0% 0.204 7.91 381.65
KNN-CE 1 100.0% 0.366 79.80 74.69 100.0% 0.225 14.08 345.56 100.0% 0.175 2.81 437.72
2 100.0% 0.380 106.05 50.75 100.0% 0.227 14.34 338.37 100.0% 0.187 4.16 418.57
4 100.0% 0.748 255.59 9.03 100.0% 0.228 16.73 326.78 100.0% 0.204 6.98 387.88
8 100.0% 1.240 467.86 3.01 100.0% 0.219 14.24 343.23 100.0% 0.199 5.27 404.10
Table 1 reports the codebook usage and reconstruction quality scores—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines. Reconstruction through softmax-smoothed quantization (i.e., without Gumbel randomness) deteriorated substantially when combined with the perplexity-based regularization (“PPL”; Equation 11). This degradation stems from the mismatch between soft quantization during the training and hard quantization at inference; the smoothed quantizers deviated from the simplex vertices, as reflected in their high individual perplexity, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$ (Table 2). These observations support the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing.
@@
$\times16$$\times$32; 1024 $\times64$$\times$3; 8192 $\times64$$\times$32; 8192
Method $K/4$ 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max
PPL 910.40 933.12 963.29 995.91 7559.89 7561.99 7563.10 7565.02 7230.60 7267.06 7315.62 7424.16
KNN-L2 1 1.17 1.59 2.36 6.26 1.00 1.00 1.00 4.00 2.24 40.45 86.53 249.64
2 1.00 1.08 1.86 27.37 1.00 1.00 1.06 4.65 1.14 1.55 2.33 7.92
4 1.00 1.06 1.86 62.99 1.00 1.00 1.00 4.61 1.00 1.06 1.80 20.42
8 1.00 1.10 1.90 120.68 3.33 3.87 5.22 12.03 1.00 1.05 1.79 107.70
KNN-CE 1 1.18 1.60 2.38 6.28 3.96 4.51 5.81 11.96 1.09 1.45 2.17 5.94
2 1.23 1.72 2.73 8.07 4.04 4.66 6.16 11.77 1.06 1.38 2.08 6.15
4 1.42 1.78 2.38 5.45 4.16 4.86 6.57 13.04 1.08 1.44 2.16 6.58
8 2.38 2.80 3.71 7.78 4.46 5.31 7.39 13.81 1.29 1.75 2.65 9.26
@@
The perplexity-based regularization was only effective when combined with the Gumbel-softmax sampling, and onehot quantization in the forward computation by STE was necessary to ensure full codebook usage (“Hard-Gumbel”). Otherwise, code collapse was not prevented when the channel dimensionality was large ($C=32$; “Soft-Gumbel”).
@@
By contrast, the proposed KNN-based regularization successfully prevented code collapse and achieved near-complete codebook utilization without resorting to Gumbel-softmax sampling. The reconstruction quality was also superior or competitive with the Gumbel-softmax + perplexity approach across all feature map sizes.
The choice of deviation metric (L2 distance vs. cross-entropy, CE) did not result in consistent global superiority, although noticeable code collapse occurred when KNN-L2 regularization was combined with hard Gumbel-softmax sampling, which nonetheless had limited effect on the reconstruction quality.
@@
In terms of computational efficiency, however, cross-entropy proved more favorable. Notably, its performance was highest when the number of neighbors was minimal ($K/4=1$) across all feature map sizes (indicated by the underlined scores in Table 1); increasing $K$ did not yield further improvements. Although most smoothed quantizers remained unregularized under this setting, they still achieved low individual perplexities after the training (Table 2), indicating a tight approximation of quantization. By contrast, minimizing only the L2 distance between the simplex vertices and their single nearest neighbor per GPU led to performance degradation for the feature map of size $64\times 64\times 8192$. Since requiring more neighbors necessitates a larger batch size (see §5.3 for details), cross-entropy emerges as a more scalable option.
Finally, both STE and rotational gradient estimation (RE) exhibited severe code collapse when the channel dimensionality was large ($C=32$), reaffirming prior findings that reducing channel dimensionality is necessary for stable training (Yu et al., 2022a, ; Yu et al.,, 2024; Mentzer et al.,, 2024). By contrast, SimVQ maintained full codebook usage across all settings and achieved strong reconstruction quality even at high dimensionality. However, the next section presents a case study in which all gradient-estimation approaches—including SimVQ—encounter code collapse, highlighting the difficulty of achieving robust prevention of code collapse across different settings.
@@ 4.2 Contrastive Learning
Table 3: Codebook usage in Wav2Vec 2.0 pretraining, evaluated on the LibriSpeech dev-clean split (similar results were observed for the other dev/test splits). The proposed method is denoted as “KNN-L2/CE”.
@@
By contrast, the proposed KNN-based regularization ensured (near-)complete code utilization in both conditions when cross-entropy was used as the divergence metric on the simplex. As in the discrete autoencoding experiments, a single neighbor per GPU was sufficient to obtain this effect. The L2 metric, on the other hand, proved insufficient to prevent code collapse, especially when combined with Gumbel sampling.
@@ Acknowledgments
Hidden for double-blind review.

### v3

@@ 2.3 Regularization
$\mathcal{L}_{\text{reg}}$ $=\mathcal{L}_{\text{hard}}$
$:=N^{-1}\sum_{i=1}^{N}\left(\beta\underbracket{\|\mathbf{z}_{i}-\mathtt{detach}(\mathbf{q}_{\iota(\mathbf{z}_{i})})\|^{2}}_{\text{Commitment Loss}}\right$.
$\qquad\qquad\qquad\left.+\underbracket{\|\mathtt{detach}(\mathbf{z}_{i})-\mathbf{q}_{\iota(\mathbf{z}_{i})}\|^{2}}_{\text{Codebook Loss}}\right)$ (10)
@@
It should be noted, however, that the codebook loss in Equation 10 does not inherently prevent code collapse. Some codebook vectors may never serve as the nearest neighbor of any pre-quantized feature and therefore receive no updates (Zhu et al.,, 2025). Accordingly, previous studies have resorted to additional workarounds; for example, unused codebook vectors may be reset to the positions of pre-quantized features (Dhariwal et al.,, 2020). More recently, Zhu et al., (2025) proposed another remedy called SimVQ, which reparameterizes the codebook $Q$ as the product of a randomly frozen matrix $Q^{\prime}\in\mathbb{R}^{M\times D}$ and a learnable matrix $W\in\mathbb{R}^{D\times D}$ (i.e., $Q=Q^{\prime}W$). In this formulation, all codebook vectors share learnable parameters with one another ($\mathbf{q}_{m}=q^{\prime}_{m,1}\mathbf{w}_{1}+\cdots+q^{\prime}_{m,D}\mathbf{w}_{D}$), so updates to one vector propagate to the others.
@@ 4 Experiments
The proposed regularization for smoothed vector quantization was benchmarked on two tasks: discrete autoencoding (§4.1) and contrastive learning (§4.2). The Python code used for these experiments will be published once this study is accepted.
@@ 4.1 Discrete Autoencoding
The first benchmark assessed the proposed regularization in the context of discrete autoencoding on the ImageNet dataset (Deng et al.,, 2009). Input images were convolutionally encoded into latent feature maps, whose pixels were subsequently quantized (Esser et al.,, 2021; Fifty et al.,, 2025). A decoder convolutional network reconstructed the input images from these quantized feature maps, and the entire model was trained to minimize the L2 reconstruction loss ($L_{\text{main}}$ in Equation 9). Further details of the network architecture and training setup are provided in Appendix B.1.
Extending prior work (Esser et al.,, 2021; Fifty et al.,, 2025), four combinations of feature map and codebook sizes were examined. In addition to the previously studied configurations of $H\times W\times C=16\times 16\times 32$ ($M=1024$) and $64\times 64\times 3$ ($M=8196$), the channel dimensionality of the latter was further increased to $C=32$ and $C=2048$, yielding feature maps of size $64\times 64\times 32$ and $64\times 64\times 2048$ (both with $M=8196$).
Each model was trained using four GPUs in parallel. Accordingly, the proposed method computed the $K/4$-nearest neighbors of each simplex vertex independently on each GPU (see §5.3 for further discussion). The number of neighbors was chosen from $K/4\in\{1,2,4,8\}$. [^3] The value of $K$ was upper-bounded at $8\times 4$ by available computational resources; With a maximum batch size of $64$, the total number of latent pixels was $64\times 64\times 64=8\times 4\times 8196$, allowing only $8\times 4$ neighbors per vertex of $\Delta^{8196}$. This implementation constraint is further discussed in §5.3. The weight $\beta$ on the commitment loss (Equation 10) for STE-based methods (including SimVQ) and rotational gradient estimation was set to 1.0, following Fifty et al., (2025) and Zhu et al., (2025).
@@
$\times16$$\times$32; 1024 $\times64$$\times$3; 8196 $\times64$$\times$32; 8196 $\times64$$\times$2048; 8196
Method $K/4$ Code Use ($\uparrow$) rMSE ($\downarrow$) FID ($\downarrow$) IS ($\uparrow$) Code Use rMSE FID IS Code Use rMSE FID IS Code Use rMSE FID IS
STE Euclid 4.5% 0.404 124.86 36.61 100.0% 0.167 7.25 402.84 1.7% 0.235 22.02 290.24 0.1% 0.274 32.60 227.73
Cosine 3.0% 0.381 117.95 41.48 70.9% 0.197 13.97 348.31 2.8% 0.186 12.11 363.86 0.2% 0.227 23.03 290.15
SimVQ 100.0% 0.340 87.26 71.33 100.0% 0.170 7.44 400.41 100.0% 0.148 3.97 436.29 70.8% 0.682 286.18 3.19
RE Euclid 3.1% 0.460 170.30 19.25 78.85% 0.171 10.21 377.56 0.5% 0.271 40.58 203.52 0.1% 0.301 53.81 153.77
Cosine 2.8% 0.423 157.77 24.33 99.5% 0.194 14.67 344.17 4.4% 0.180 10.83 372.89 1.3% 0.193 16.98 331.17
HG PPL 100.0% 0.349 100.29 54.04 100.0% 0.189 19.07 321.16 100.0% 0.163 10.33 384.04 100.0% 0.165 8.89 392.77
KNN-CE 1 100.0% 0.368 86.49 68.27 100.0% 0.226 14.27 342.64 100.0% 0.173 3.17 435.04 100.0% 0.177 3.42 432.08
SG PPL 55.0% 0.826 173.59 9.61 100.0% 0.183 10.28 377.79 48.4% 0.386 45.61 167.14 34.6% 1.253 368.71 1.58
KNN-CE 1 100.0% 0.569 194.74 16.60 100.0% 0.233 15.41 329.78 100.0% 0.186 3.83 425.33 100.0% 0.188 3.73 422.84
Softmax PPL 100.0% 1.112 309.60 4.55 81.6% 0.701 37.99 196.72 99.8% 0.725 83.07 77.98 89.3% 1.345 84.52 80.39
KNN-L2 1 100.0% 0.358 76.29 81.61 100.0% 0.224 12.08 354.36 100.0% 0.387 32.20 210.61 75.9% 1.259 104.22 56.87
2 100.0% 0.366 81.69 71.75 100.0% 0.226 12.04 348.48 100.0% 0.175 3.62 427.42 58.3% 1.302 94.82 62.82
4 100.0% 0.379 87.05 62.37 100.0% 0.205 9.01 383.56 100.0% 0.196 5.64 403.06 68.3% 1.283 96.45 73.56
8 99.9% 0.343 73.72 88.53 100.0% 0.199 12.64 361.23 100.0% 0.204 7.91 381.65 80.8% 1.134 73.79 92.46
KNN-CE 1 100.0% 0.366 79.80 74.69 100.0% 0.225 14.08 345.56 100.0% 0.175 2.81 437.72 100.0% 0.424 18.58 289.42
2 100.0% 0.380 106.05 50.75 100.0% 0.227 14.34 338.37 100.0% 0.187 4.16 418.57 100.0% 0.212 4.22 416.56
4 100.0% 0.748 255.59 9.03 100.0% 0.228 16.73 326.78 100.0% 0.204 6.98 387.88 100.0% 0.190 4.70 411.70
8 100.0% 1.240 467.86 3.01 100.0% 0.219 14.24 343.23 100.0% 0.199 5.27 404.10 100.0% 0.201 3.77 416.71
Table 1 reports the codebook utilization and reconstruction performance—including root mean squared error (rMSE), Inception Score (IS; Salimans et al.,, 2016), and Fréchet Inception Distance (FID; Heusel et al.,, 2017)–for the proposed method (“KNN-L2/CE”) and the baselines. Reconstruction quality deteriorated substantially when softmax-smoothed quantization (i.e., without Gumbel randomness) was combined with the perplexity-based regularization (“PPL”; Equation 11). This degradation arises from a mismatch between soft quantization during the training and hard quantization at inference; the smoothed quantizers deviated from the simplex vertices, as reflected by their high individual perplexity, $\exp(-\sum_{m=1}^{M}p_{m}\log p_{m})$ (Table 2). These results support the argument made in the Introduction that perplexity-based regularization alone does not promote tight smoothing.
@@
$\times16$$\times$32; 1024 $\times64$$\times$3; 8192 $\times64$$\times$32; 8192 $\times64$$\times$2048; 8192
Method $K/4$ 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max 75% 90% 99% Max
PPL 910.40 933.12 963.29 995.91 7559.89 7561.99 7563.10 7565.02 7230.60 7267.06 7315.62 7424.16 8191.71 8191.85 8192.02 8192.61
KNN-L2 1 1.17 1.59 2.36 6.26 1.00 1.00 1.00 4.00 2.24 40.45 86.53 249.64 8191.66 8191.82 8191.98 8192.34
2 1.00 1.08 1.86 27.37 1.00 1.00 1.06 4.65 1.14 1.55 2.33 7.92 8191.49 8191.72 8191.94 8192.35
4 1.00 1.06 1.86 62.99 1.00 1.00 1.00 4.61 1.00 1.06 1.80 20.42 8191.63 8191.80 8191.97 8192.36
8 1.00 1.10 1.90 120.68 3.33 3.87 5.22 12.03 1.00 1.05 1.79 107.70 8191.69 8191.82 8191.97 8192.41
KNN-CE 1 1.18 1.60 2.38 6.28 3.96 4.51 5.81 11.96 1.09 1.45 2.17 5.94 1.90 124.24 503.28 1024.38
2 1.23 1.72 2.73 8.07 4.04 4.66 6.16 11.77 1.06 1.38 2.08 6.15 1.28 2.24 16.83 676.31
4 1.42 1.78 2.38 5.45 4.16 4.86 6.57 13.04 1.08 1.44 2.16 6.58 1.36 1.98 3.16 11.94
8 2.38 2.80 3.71 7.78 4.46 5.31 7.39 13.81 1.29 1.75 2.65 9.26 1.86 2.51 4.28 26.69
Figure 2: (Top) Histogram of per-image codebook utilization on the validation dataset, defined as the percentage of codes used at least once within each image. (Bottom) Histogram of code popularity, measured as the proportion of validation images that contain each code.
@@
Perplexity-based regularization was effective only when combined with the Gumbel-softmax sampling, and onehot quantization in the forward pass via STE was necessary to ensure full codebook utilization (“Hard-Gumbel”). Without hard quantization, code collapse was not prevented at higher channel dimensionalities ($C\geq 32$; “Soft-Gumbel”).
@@
By contrast, the proposed KNN-based regularization consistently prevented code collapse and achieved near-complete codebook utilization without resorting to Gumbel-softmax sampling. Reconstruction performance was also superior to, or competitive with, the Gumbel-softmax plus perplexity baseline across all feature map sizes.
The choice of deviation metric (L2 distance vs. cross-entropy, CE) did not exhibit a consistent global advantage. One notable exception occurred at extremely high channel dimensionality ($C=2048$). In this regime, KNN-L2 failed to achieve full codebook utilization, whereas KNN-CE maintained complete utilization. Moreover, the soft quantizers under KNN-L2 failed to concentrate around the simplex vertices in the high-dimensionality condition, as indicated by their near-maximal individual perplexity values (Table 2). This looseness in smoothing also led to degraded reconstruction quality (Table 1).
@@
Cross-entropy was additionally favorable in terms of computational efficiency. Its performance was highest when using the minimum number of neighbors ($K/4=1$) for nearly all feature map sizes (underlined entries in Table 1); increasing $K$ yielded no further gains except in the extreme case of $C=2048$. Although most soft quantizers remained formally unregularized under this setting, they nonetheless achieved low individual perplexity after training (Table 2), indicating a tight approximation to hard quantization. By contrast, minimizing the L2 distance to only a single nearest neighbor per GPU led to performance degradation for feature maps of size $64\times 64\times\{32,2048\}$. Because increasing the number of neighbors requires a larger batch size (see §5.3 for details), cross-entropy emerges as the more scalable option.
Finally, both STE and rotational gradient estimation (RE) exhibited severe code collapse at higher channel dimensionalities ($C\geq 32$), reaffirming prior findings that reducing channel dimensionality is critical for stable training (Yu et al., 2022a, ; Yu et al.,, 2024; Mentzer et al.,, 2024). SimVQ was more robust, maintaining full codebook utilization and strong reconstruction performance at $C=32$. However, when the channel dimensionality increased to $C=2048$, its performance fell below that of classical STE-based models. Although SimVQ still utilized 70.8% of the codebook overall, each individual image was encoded by using no more than 1% of the codes (Figure 2). Effective vector quantization instead requires diverse code usage within individual images (typically 10–40%), a property that KNN-CE regularization preserves across channel dimensionalities ($C\in\{3,32,2048\}$).
The next section presents a case study in which all gradient-estimation approaches—including SimVQ—exhibit more pronounced code collapse, highlighting the difficulty of achieving robust prevention of code collapse across different settings.
@@ 4.2 Contrastive Learning
Table 3: Codebook usage in Wav2Vec 2.0 pretraining, evaluated on the LibriSpeech dev-clean split (similar results were observed for the other dev/test splits).
@@
By contrast, the proposed KNN-based regularization ensured high code utilization in both conditions. As in the discrete autoencoding experiments, a single neighbor per GPU was sufficient to obtain this effect.
@@ Acknowledgments
This study was supported by JSPS Grant-in-Aid for Scientific Research A (JP24H00774), B (JP22H03914), and C (JP24K15087); JST AIP Accelerated Program (JPMJCR25U6) and Core Research for Evolutional Science and Technology (JPMJCR22P5); and Kayamori Foundation of Informational Science Advancement (K35XXVIII620). The author also gratefully acknowledges the support of the Academic Center for Computing and Media Studies, Kyoto University, regarding the use of their supercomputer system.